OpenStudy (anonymous):

Divide and express the result in standard form: 9i ----- 6-8i

OpenStudy (anonymous):

multiply top and bottom by the conjugate of the denominator. the conjugate of \(a+bi\) is \(a-bi\) this works because \(a+bi)(a-bi)=a^2+b^2\) a real number start with \[\frac{9i}{6-8i}=\frac{9i}{6-8i}\times \frac{6+8i}{6+8i}\]

OpenStudy (anonymous):

\[(36-27i)/-14\]

OpenStudy (anonymous):

the denominator will be \(6^2+8^2=100\) and the numerator will be whatever you get when you multiply

OpenStudy (anonymous):

be careful, because it is not the answer given above